OpenStudy (anonymous):

Multiply the equation by a power of 10 to write an equivalent equation with integer coefficients. 6.2x + 4.5 = 3.8x + 7.9

OpenStudy (mathstudent55):

Since every coefficient in the equation has a decimal number that goes up to the tenth place, you need to multiply the entire equation by 10 to make all coefficients integers. 10(6.2x + 4.5) = 10(3.8x + 7.9) Distribute the 10 on both sides: 10(6.2x) + 10(4.5) = 10(3.8x) + 10(7.9) Do all those multiplications by 10: 62x + ... Can you continue?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

62x + 45 = 38x +79

OpenStudy (anonymous):

is that right?

OpenStudy (mathstudent55):

Correct.

OpenStudy (anonymous):

Should I simplify it anymore?

OpenStudy (mathstudent55):

You have doen what the problem asks. There is no factor you can take out because 62, 45, 38, and 79 have no common factor.